# On the foundation of quantum decision theory

### v3

@@ Abstract
Quantum decision theory is introduced here, and a new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, and next on arguments for the Born rule, which is the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that each accessible variable can be seen as a function of a specific inaccessible variable. Another assumption is that there exist two maximal accessible decision processes in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example.
Keywords: Accessible decision variable; action; Born rule; decision; Hilbert space formalism; illustrating example; quantum probabilities.
@@ 1 Introduction
Political decisions are important, and may have large consequences. In an autocratic country political decisions are made by sovereign leaders, who have gained their positions through power exchanges. In democratic countries, at least in principle people may control their leaders’ decisions.
In this article we will concentrate on decisions made by single persons. From a psychological point of view, such decisions have been thoroughly reviewed by Newell et al. (2015). The book emphasizes the relationship between learning and decisions, arguing that the best way to understand how and why decisions are made is in the context of learning and knowledge acquisition that precedes them, and the feedback that follows. Some of the discussions in op. cit. are also relevant for the developments in the present article.
@@
A person always has his history, and through his history he learns how to handle decisions. Hence learning is important. In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory.
Admittedly, the decision foundation based upon optimizing expected utility has had and has still strong theoretical support. There are several axiom systems that can motivate such a basis, and the mathematics behind such axioms can be traced back to Bernoulli and Pascal. Newell et al. (2015) discuss one such axiom, the sure thing principle: If a person prefers option A to option B under the condition X, and he also prefers A to B under the opposite condition not-X, then he always prefers A to B.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory. This theory is completely based upon quantum mechanics, whose foundation is discussed in the book Helland (2021) and in some newer articles Helland (2022a,b, 2023a,b).
Pionering articles on quantum decision theory are Aerts (2009) and Yukalov and Sornette (2014); see also the book by Busemeyer and Bruza (2012). More recent review articles are Ashtiani and Azgomi (2015) and Pothos an Busemeyer (2022). Our approach will be independent of this literature; a completely new entrance to the theory will be aimed at.
The plan of this article is as follows: I first review briefly a new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications are then given, and used to illustrate some of the peculiarities of quantum probabilities. Then the interpretation is briefly discussed before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b, 2023a,b,c,2024a). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,...$ are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
@@
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take ϕ=italic-$\phi=$ (position, momentum) and ϕ=italic-$\phi=$ full spin vector. In the last example, in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$. Here these variables are theoretical, variables coupled to some physical situation. Following Zwirn (2016, 2020), every description of reality must be seen from the point of view of some observer. Hence we can assume that the variables also exist relative to $C$.
@@
But observers may communicate. The mathematical model developed in the articles above is equally valid relative to a group of people who can communicate about physics and the various theoretical variables. This gives a new version of the theory, a version where all theoretical variables are defined jointly for such a group. The only difference here is that, for the variables to function during the communication, it must always be possible to define them in words.
@@
In the above two examples there are also maximal accessible variables: In the first example either position or momentum, in the second example the spin component $\theta^{a}$ in some direction $a$. In general, an accessible variable $\theta$ is maximal if there is no other accessible variable $\lambda$ such that $\theta=f(\lambda)$ for some non-invertible function $f$.
@@
Two different accessible variables $\theta$ and $\eta$ are said to be related if there is a transformation $k$ in $\phi$-space and a function $f$ such that $\theta=f(\phi)$ and $\eta=f(k\phi)$. In Helland (2023b) it is shown that if both $\theta$ and $\eta$ are maximal and take $r$ values, then $\phi$ can be constructed in a natural way, there is an accessible variable $\lambda$ that is in one-to-one correspondence with $\theta$, all this such that $\lambda$ and $\eta$ are related according to this definition.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ when they are different maximal accessible variables. Such variables are called - following Niels Bohr - for complementary.
@@
In the present paper, we will concentrate on accessible variables that take a finite number of values, say $r$. Then the following first theorem is proved in the above papers: Given a situation with two different maximal accessible variables $\theta$ and $\eta$, there corresponds to every accessible variable a self-adjoint operator in some complex Hilbert space $\mathcal{H}$. With ‘different’ I here mean really different, not two maximal variables that are one-to-one functions of each other. And a ‘self-adjoint operator’ is just a complex $r\times r$ matrix with real eigenvalues. In particular then, there is such an operator $A^{\theta}$ corresponding to $\theta$ and an operator $A^{\eta}$ corresponding to $\eta$.
@@
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$.
If $\theta$ is maximal as an accessible variable, the eigenvectors of $A^{\theta}$ have a similar interpretation. In my opinion, such eigenvectors, where $\theta$ is some meaningful variable, should be taken as the only possible state vectors. These have straightforward interpretations, and from this version of the theory also a number of so-called ‘quantum paradoxes’ can be illuminated, see Helland (2022b,d).
This not only points at a new foundation of quantum theory, and it also suggests a general epistemic interpretation of the theory: Quantum theory is not directly a theory about the world, but a theory about an actor’s knowledge of the world. Versions of such an interpretation already exist, and it is one of the very many suggested interpretations of quantum mechanics.
@@
A ket vector $|\psi\rangle$, a state vector, can be seen as a unit column vector in $\mathcal{H}$. I will concentrate on state vectors of the form $|\psi\rangle=|b;j\rangle$, an eigenvector of an operator $A^{b}$, in my language, equivalent to a particular answer to a question of the form ‘What is $\theta^{b}$?’, where $\theta^{b}$ is an accessible theoretical variable.
To any ket vector $|a;k\rangle$ there corresponds a bra vector $\langle a;k|$, the corresponding complex conjugate row vector. Scalar products between two state vectors can then be written $\langle a;k|b;j\rangle$. I will also need scalar products of the form $\langle a;k|A|a;k\rangle$, where $A$ is an operator.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values $1,...,r$, such that $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$. If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in correspondence with a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make up his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to make a decision. Let two different maximal, decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$, and $\eta=j$ corresponds to the action $b_{j}$ $(j=1,...,r)$. Then, by the theory of the previous subsection, we can model the situation by using quantum theory.
Note that this simple model does not cover all situations. Sometimes we have a choice between an infinite number of possibilities, and sometimes the outer context changes during the decision process. Nevertheless, the simple model is a good starting point.
@@ 2.3 The Hilbert space formalism for decisions
It is well known that our minds may be limited, for instance when faced with difficult decisions. We will first mention a side result in this direction from the present development.
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneously have in mind any other maximal accessible variable which is related to $\theta$, but not related to $\eta$. It was claimed by Helland (2022c) that the violation of a famous inequality by practical Bell experiments can be understood on the basis of this theorem. See also Helland (2023c).
Note that this result has the qualification ‘at the same time’, and indicates a specific restriction to two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
For the present article, however, the direct results from Helland (2022b, 2023b) are more important. Consider again a decision situation, and assume the simple model of the previous Subsection. In particular let $C$ at the same time be confronted with at least two different maximal related decision processes. Then the following hold:
- Each decision variable $\theta$ is associated with a self-adjoint operator $A^{\theta}$, whose eigenvalues are that possible values of $\theta$.
@@
- In the maximal case, the eigenvectors of the operator can be given interpretation: They are coupled to one particular decision process and a specific choice in this decision process: In concrete terms, the eigenvectors $v$ are in one-to-one correspondence with 1) some maximal accessible decision variable $\theta$, and 2) a specific value $u$ of $\theta$. In other words, the possible eigenvectors are in one-to-one correspondence with 1) the question ‘Which decision process?’ and 2) ‘Which action did this decision process lead to?’.
@@
This can be taken as a starting point of quantum decision theory, but to develop this theory further, we need to be able to calculate probabilities of the various decisions.

### v4

# A new foundation of quantum decision theory

@@ Abstract
Quantum decision theory is introduced here, and new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, next on arguments for the Born rule, that is, the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that all accessible variables can be seen as functions of a specific inaccessible variable. Another assumption is that there exist two maximal accessible theoretical variables in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example. Finally, a broad discussion of decision processes is given.
Keywords: Accessible decision variabel, action, Born rule; decision, Hilbert space formalism, quantum probabilities.
@@ 1 Introduction
Political decisions are important, and may have large consequences. In an autocratic country political decisions are made by sovereign leaders, having gained their positions through power exchanges. In democratic countries, at least in principle people may control their leaders’ decisions.
In this article I will concentrate on decisions made by single persons. From a psychological point of view, such decisions have been thoroughly reviewed by Newell et al. (2015). The book emphasizes the relationship between learning and decisions, arguing that the best way to understand how and why decisions are made is in the context of learning and knowledge acquisition which precedes them, and the feedback which follows. Some of the discussions in op. cit. are also relevant for the developments in the present article.
@@
A person always has his history, and through his history he learns how to handle decisions. Hence learning is important. A person’s learning history starts when he or she is a child, when his or her ideals usually are the parents. Later, he or she is influenced by other people, but we can never disregard such influences as a background for a person’s decisions, even in cases where the person claims that his aim is just to optimize his behaviour according to some criterion.
In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory. The argument will be further elucidated later in this article.
Admittedly, the decision foundation based upon optimizing expected utility has had and has still a strong theoretical support. There are several axiom systems that can motivate such a basis, and the mathematics behind such axioms can be traced back to Bernoulli and Pascal. Newell et al. (2015) discuss one such axiom, the sure thing principle: If a person prefers option A to option B under the condition X, and he also prefers A to B under the opposite condition not-X, then he always prefers A to B.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory (QDT). This theory is completely based upon quantum mechanics, whose foundation is discussed in the book Helland (2021) and in some newer articles, Helland (2022a,b,c, 2023a,b,c, 2024a,b,c).
Pionering articles on quantum decision theory are Aerts (2009) and Yukalov and Sornette (2014); see also the book by Busemeyer and Bruza (2012). More recent review articles are Ashtiani and Azgomi (2015) and Pothos an Busemeyer (2022). A motivation and more references can also be found in the article by Wang et al. (2013). My approach will be independent of this literature; a completely new entrance to the theory will be aimed at.
The plan of this article is as follows: I first review briefly a new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications is then given, and used to illustrate some of the peculiarities of quantum probabilities. Then my foundation of QDT is further discussed from several points of view before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b,c 2023a,b,c,2024a,b,c). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,...$ are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
@@
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take ϕ=italic-$\phi=$ (position, momentum) and ϕ=italic-$\phi=$ full spin vector. In the last example, say in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$.
Here these variables are theoretical variables coupled to some physical situation. If necessary, an observed variable can be modeled as a theoretical variable plus some random error. Following Zwirn (2016, 2020), every description of reality must be seen from the point of view of some observer. Hence we can assume that the variables also exist relative to $C$.
@@
But observers may communicate. The mathematical model developed in the articles above is equally valid relative to a group of people that can communicate about the physics and about the various theoretical variables. This gives a new version of the theory, a version where all theoretical variables are defined jointly for such a group. The only difference here is that, for the variables to function during the communication, they must always be possible to define them in words.
@@
In the above two examples there are also maximal accessible variables: In the first example either position or momentum, in the second example the spin component $\theta^{a}$ in some direction $a$. From a mathematical point of view, an accessible variable $\theta$ is called maximal if there is no other accessible variable $\lambda$ such that $\theta=f(\lambda)$ for some non-invertible function $f$. In general, the term ‘maximal’ will then be seen to be maximal with respect to the partial ordering of variables given by $\alpha\leq\beta$ iff $\alpha=f(\beta)$ for some function $f$.
@@
Two different acccessible variables $\theta$ and $\eta$ are said to be related if there is a transformation $k$ in $\phi$-space and a function $f$ such that $\theta=f(\phi)$ and $\eta=f(k\phi)$. In Helland (2024a) it is shown that if both $\theta$ and $\eta$ are maximal and take $r$ values, then $\phi$ can be constructed in a natural way such that there is an accessible variable $\lambda$ that is in one-to-one correspondence with $\theta$, and such that $\lambda$ and $\eta$ are related according to this definition.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ even when they are different maximal accessible variables. Variables that are maximal, different and not in one-to-one correspondence, are called - following Niels Bohr - for complementary.
@@
In the present paper, I will concentrate on accessible variables that take a finite number of values, say $u_{1},u_{2},...,u_{r}$. Then it can be shown that pairs of maximal accessible variables are either related, are in one-to-one correspondence or have a combination of these two properties. The following Theorem is proved in Helland (2024a):
Theorem 1 Assume that there exist two different maximal accessible variables $\theta$ and $\eta$, each taking $r$ values, and not in one-to-one correpondence. Then, there exists an $r$-dimensional Hilbert space $\mathcal{H}$ describing the situation, and every accessible variable in this situation will have an associated self-adjoint operator in $\mathcal{H}$.
@@
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues of the corresponding operator are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$, where $u$ is an eigenvalue of $A^{\theta}$.
If $\theta$ is maximal as an accessible variable, the eigenvectors of $A^{\theta}$ have a similar interpretation. In my opinion, such eigenvectors, where $\theta$ is some meaningful variable, should be tried to be taken as the only possible state vectors. These have straightforward interpretations, and from this version of the theory, also a number of so-called ‘quantum paradoxes’ can be illuminated, see Helland (2023a,c).
This not only points at a new foundation of quantum theory, and it also suggests a general epistemic interpretation of the theory: Quantum theory is not directly a theory about the world, but a theory about an actor’s knowledge of the world. Versions of such an interpretation already exist, and they are among the very many suggested interpretations of quantum mechanics.
@@
A ket vector $|\psi\rangle$, a state vector, can be seen as a unit column vector in $\mathcal{H}$. I will concentrate mainly on state vectors of the form $|\psi\rangle=|b;j\rangle$, an eigenvector of an operator $A^{b}$, in my language, equivalent to a particular answer to a question of the form ‘What is $\theta^{b}$?’, where $\theta^{b}$ is a maximal accessible theoretical variable.
To any ket vektor $|a;k\rangle$ there corresponds a bra vector $\langle a;k|$, the corresponding complex conjugate row vector. Scalar products between two state vectors can then be written $\langle a;k|b;j\rangle$. I will also need scalar products of the form $\langle a;k|A|a;k\rangle$, where $A$ is an operator.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he or she has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values $1,...,r$, such that $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$. If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in relation to a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to take a decision. Let two different maximal decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$, and $\eta=j$ corresponds to the action $b_{j}$ $(j=1,...,r)$. Then, by the theory of the previous subsection, we can model the situation by using quantum theory.
Note that both $\theta$ and $\eta$ may be vector variables. Say that $\theta=(\theta_{1},...,\theta_{n})$, where each $\theta_{j}$ is a simpler decision variable. Then this corresponds to a situation where $C$, in addition to the difficult decision given by $\eta$, is faced with $n$ more simple decisions. Such a situation is not uncommon. In each situation where we shall make a difficult decision, we will be in a context where also a number of trivial decisions may have to be made just in order to survive and to function well in the given context. For many people, these trivial decision occupy a large portion of their mind, such a large part that the vector decision variable $\theta$ also must be considered to be maximal. The assumption that both $\theta$ and $\eta$ take the same number of values $r$, can be satisfied by artificially adding some actions to one of the decision problems.
The simple model above does not cover all situations. Sometimes we have a choice between an infinite number of possibilities, and sometimes the outer context changes during the decision process. Nevertheless, the simple model is a good starting point.
Another simple model is the following: Let $\eta$ be a maximal accessible of some kind, describing the relevant state of the observer $C$, and let $\theta$ be a maximal accessible decision variable. Theorem 1 is equally valid in this situation.
In this article, I will mainly focus on maximal decision problems. A thorough discussion including decisions that are not maximal, is given in Mogiliansky et al. (2009).
2.3 When is a macroscipic accessible variable maximal?
For the purpose of this article, it is crucial that my foundation of quantum theory also is valid when relevant accessible variables are macroscopic. In the language of Khennikov (2010), I also want to include quantum-like models, which are applicable in biology, economy, psychology and many other disciplines. Quantum structure may be ubiquitous. A special case is quantum cognition theory (Wang et al., 2013; Pothos and Busemeyer, 2022), which includes quantum decision theory.
The crucial assumption of Theorem 1 is that there exist two different maximal accessible theoretical variables in some given situation. This situation may well be macroscopic.
I will start with an example from Helland (2021). In a medical experiment, let $\mu_{a},\mu_{b},\mu_{c}$ and $\mu_{d}$ be continuous inaccessible parameters, the hypothetical effects of treatment $a,b,c$ and $d$, respectively. Assume that the focus of the experiment is to compare the effect of treatment $b$ with the mean effect of the other treatments, which is supposed to give the parameter $\frac{1}{3}(\mu_{a}+\mu_{c}+\mu_{d})$. One wants to do a pairwise experiment, but it turns out that the maximal parameter that can be estimated is $\theta^{b}=\mathrm{sign}(\mu_{b}-\frac{1}{3}(\mu_{a}+\mu_{c}+\mu_{d}))$.
In the language of the present article, $\theta^{b}$ here is a maximal accessiblble theoretical variable. In concrete terms, one can imagine the following situation: One has four different ointments against rash. A patient is treated with ointment $b$ on one side of his back, a mixture of the other ointments on the other side of his back. It is only possible to observe which side improves best but this observation is assumed to be very accurate. In practice the experiment is done on many patients, and one selects out the patients were the difference is clear.
The concrete factors above make $\theta^{b}$ maximal. For a complete description of the experiment, see Helland (2021), Subsection 5.6.4. The application of a quantum-like model to the important field of experimental design is a completely new area that ought to be treated more thoroughly.
Several examples are given in Haven and Khrennikov (2013), say, and similar discussions could have been made with some of these examples. I will maintain, however, that the term maximal has a precise definition, as given in Subsection 2.1. Let first a partial ordering among theoretical variables be defined by $\alpha\leq\beta$ if $\alpha=f(\beta)$ for some function $f$. An accessible variable $\theta$ is maximal if it is maximal among the accessible variables with respect to this partial ordering.
In a setting with decision variables, such a variable can be seen to be maximal if the corresponding actor(s) just is/are able to make the decision.
@@ 2.4 The Hilbert space formalism for decisions
It is well known that our minds may be limited, for instance when faced with difficult decisions. I will first mention a side result in this direction from the present development.
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneous have in mind any other maximal accessible varable which is related to $\theta$, but not related to $\eta$. It was claimed in Helland (2022c) that the violation of a famous inequality by practical Bell experiments, can be understood on the basis of this theorem. See also Helland (2023b), where a corresponding theorem is formulated without any symmetry assumption.
Note that this result has the qualification ‘at the same time’, and indicate a specific restriction to two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
For the present article, however, the direct results from Helland (2022b, 2024a) are equally important. Consider again a decision situation, and assume the simple model of the previous Subsection. In particular let $C$ at the same time be confronted with at least two different maximal related decision processes. Then the following hold:
- Each decision variable $\eta$ is associated with a self-adjoint operator $A^{\eta}$, whose eigenvalues are the possible values of $\eta$.
@@
- In the maximal case, the eigenvectors of the operator can be given interpretations: They are coupled to one particular decision process and a specific choice in this decision process: In concrete terms, the eigenvectors $v$ are in one-to-one correspondence with 1) some maximal accessible decision variable $\eta$, and 2) a specific value $u$ of $\eta$. In other words, the possible eigenvectors are in one-to-one correspondence with 1) the question ‘Which decision process?’ and 2) ‘Which action did this decision process lead to?’.
@@
This can be taken as a starting point of quantum decision theory, but to develop this theory further, we need to be able to calculate probabilities for the various decisions.